Ticket Number: HEA-NOC/20030802-2 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: Eircom
Ticket Opened: 20030802 12:44 UTC+1 Problem Start: 20030802 04:51 UTC+1
Ticket Closed: 20031202 10:02 UTC Problem End: 20031119 11:48 UTC
Site/line: Cork - Limerick STM-1 (DPL 560011), Cork - Limerick STM-1
(DPL 560205), Cork PoP, Limerick PoP
There have been an excessive number of SONET alarms on the Cork-Limerick
backup Circuit since 04:51
Circuit has remained up and this problem is not service-affecting.
20030802 12:25 CMC Contacted Eircom Primary and secondary POC.
20030802 18:34 CMC Eircom advise they have performed checks at the SDH
level which show an all-clear, issue is covered by ticket id #2114 and
will remain open until at least tuesday.
20030804 16:40 CMC Burst of EONET alarms observed on the primary
Cork-Limerick curcuit, however issue has not proved service-affecting thus
20030806 15:08 CS Further alarms observed on both the backup and primary
Cork-Limerick circuits, still not service affecting. Eircom have left the
ticket id #2114 open as a result and will perform more checks.
20030807 13:10 CW Vanessa called to infor that SDH was clear at 0800 this
20030808 14:30 CS Contacted Eircom to inform of more SONET alarms on both
circuits. Issued ticket #524 for primary circuit which wasn't reported
20030808 14:50 CS Was contacted by Sean in Eircom, no errors detected at
SDH level, intrusive tests will need to be scheduled. Ticket #2114 for
backup circuit has been closed and re-opened as #524 due to Eircom
internal ticketing procedures.
20030813 10:43 CS Intrusive testing at Cork & Limerick POPs has been
scheduled to begin at 8am Thursday (20030814).
20030818 15:01 CS Intrusive testing has been re-scheduled to begin at 0830
Wednesday (20030820) as previous testing did not take place. There have
been more SONET alarms since last ticket update.
20030820 10:33 CS Intrusive tests took place, no errors or problems were
found. Will continue to monitor for alarms.
20030820 11:38 JL
Eircom rang to say they can find no errors on the line.Fibre was cleaned
and intrusive tests run.
20031202 10:00 AH Card replaced in Limerick as per Ticket #20031118-3.
Circuit now operating in a backup role only.
This ticket can be monitored at http://www.hea.net/tickets/20030802-2